find bounded regoin y=X2 AND y4

Grade:10

1 Answers

Sher Mohammad IIT Delhi
10 years ago
y=x^2
y^4=x^8
y=y^4 gives point as x=0, x=1 , x=-1 ,

Sher Mohammad
B.Tech, IIT Delhi